Let T: R^2→R^2 be a linear opeator with matrix


[ 7 1]


[-1 1] (w.r.t. the standard basis). Use Cayley haMilton theorem to check whether T is invertible or not. If T is invertible, obtain T^-1(x,y) for (x,y)∈ R^2. If T is not invertible, obtain the minimal polynomial of T.


1
Expert's answer
2021-02-28T17:19:28-0500

Solution:

Given matrix, "A=\begin{bmatrix} 7 \ \ 1\\-1\ 1\end{bmatrix}"

First we find characteristic polynomial, "p(t)=\begin{vmatrix} A-tI\end{vmatrix}"

"=\begin{vmatrix} 7-t \ \ \ \ 1\\-1\ \ \ 1-t\end{vmatrix}"

"=(7-t)(1-t)-1(-1)
\\ =7-t-7t+t^2+1
\\ =t^2-8t+8"

Now, applying Cayley hamilton theorem,

"O=p(A)=A^2-8A+8I"

where "O" is zero matrix and "I" is identity matrix of order 2.

So, "A^2-8A=-8I"

"\Rightarrow-\dfrac18A^2+A=I"

"\Rightarrow A(-\dfrac18A+I)=I" ...(i)

Similarly, we can write this as: "(-\dfrac18A+I)A=I" ...(ii)

Thus, from (i) and (ii), matrix "(-\dfrac18A+I)" is the inverse of matrix "A" .

"\therefore A^{-1}=(-\dfrac18A+I)"

"=-\frac18\begin{bmatrix} 7 \ \ 1\\-1\ 1\end{bmatrix}+\begin{bmatrix} 1 \ \ 0\\0\ 1\end{bmatrix}"

"=\begin{bmatrix} -\frac78 \ \ -\frac18\\\frac18\ -\frac18\end{bmatrix}+\begin{bmatrix} 1 \ \ 0\\0\ 1\end{bmatrix}"

"=\begin{bmatrix} \frac18 \frac{-1}8\\\frac18\ \frac78\end{bmatrix}"

This is required answer.